# Mpox-AISM : AI-Mediated Super Monitoring for Mpox and Like-Mpox

### v3

@@ Abstract
The key to preventing the spread of mpox (monkeypox) lies in timely, convenient, and accurate diagnosis for earlier-stage infected individuals. Unfortunately, the resemblances between common skin diseases and mpox and the need for professional diagnosis inevitably deteriorated the diagnosis of earlier-stage patients with Mpox and contributed to its widespread outbreak in crowded areas. Here, we proposed a real-time visualization strategy called "Super Monitoring" using artificial intelligence and Internet technology, thereby performing a low-cost, convenient, timely, and unspecialized diagnosis for earlier-stage mpox. Specifically, such AI-mediated "super monitoring" (Mpox-AISM) invokes a framework assembled by deep learning models, data augmentation, self-supervised learning, and cloud services. Verified by publicly available datasets, the Precision, Recall, Specificity, and F1-score of Mpox-AISM in diagnosing mpox achieved 99.3%, 94.1%, 99.9%, and 96.6%, respectively. Furthermore, Mpox-AISM’s overall accuracy reaches 94.51% in diagnosing mpox, six like-mpox skin diseases, and normal skin. We also employed gradient-weighted class activation mapping to explain the decision-making process of Mpox-AISM, thus handily understanding the specific characteristics that may indicate the mpox’s onset and improving its reliability. With the help of the Internet and communication terminal, Mpox-AISM can perform a real-time, low-cost, and convenient diagnosis for earlier-stage mpox in various real-world settings, thereby effectively curbing the spread of mpox virus.
@@ 1 Introduction
Given the above issues, artificial intelligence (AI) offers a reliable solution [10]. Recently, deep learning (DL) models in AI have achieved significant success in various fields, including machine vision, medical imaging, and driverless vehicles. They have become sophisticated tools for addressing previously unknown and challenging problems. Moreover, AI-based innovations have reduced the cost of expensive equipment and professional knowledge required in the clinical diagnosis process, thereby helping overcome resource-constrained environments [11]. In this work, we employed the Internet, popular DL algorithms, data augmentation and a novel self-supervised learning (SSL) strategy, i.e. "A Simple Framework for Contrastive Learning of Visual Representations" (SimCLR) [12], to design a real-time and online strategy for diagnosing mpox. Our strategy considers the clinical characteristics of mpox rash and its high similarity to common skin diseases and helps diagnose earlier-stage mpox conveniently and swiftly. We validated the effectiveness of Mpox-AISM using publicly available datasets. Furthermore, we classified the mpox rash images into four grades (I, II, III, and Others) and two stages (earlier and later) based on dataset characteristics and the clinical evolution trend of the mpox rash. In detail, Grade-I means body parts are more prone to pox rash, i.e., the faces, necks, and hands [13, 14]. At the same time, Grade-II contains the arms and legs, which are not easily covered by clothing and other articles of daily use. We retested Mpox-AISM and found that it still performs well in diagnosing mpox images in different parts and stages. Particularly, to improve the reliability and security of Mpox-AISM, we also visualized and explained its decision-making process and results.
@@ 4 Limitations of study
AICOM-MP [36] 2024 SL
6124 labeled images
mpox, non-mpox
× 96.3% ×
MobileNetV3-large [37] 2024 SL
400 labeled images
mpox, chickenpox, acne, normal
× 88.2% ×
PoxNet22[38] 2023 SL
3192 labeled images
mpox, non-mpox
× 100.0% ×
DenseNet201[39] 2023 SL
1710 labeled images
mpox, chickenpox, measles, normal
× 97.6% ×
MobileNetV3-Small[40] 2023 SL
2056 labeled images
mpox, non-mpox
× 96.0% ×
ResNet18 [41] 2023 SL
3192 labeled images
mpox, non-mpox
× 99.5% ×
Vision Transformer [42] 2023 SL
3192 labeled images
mpox, non-mpox
× 94.7% ×
MonkeyNet [43] 2023 SL
8689 labeled images
mpox, chickenpox, measles,normal
× 93.2% ×
Xception+Denseet169 [44] 2022 SL
1754 labeled images
mpox, chickenpox, measles, normal
× 87.13% ×
MobileNetV2 [45] 2022 SL
3192 labeled images
mpox, non-mpox
× 91.1%
@@ 5 STAR METHODS
ResNet He et al. [53] https://github.com/pytorch/vision/blob/main/torchvision/models/resnet.py

### v4

@@ Abstract
Swift and accurate diagnosis for earlier-stage monkeypox (mpox) patients is crucial to avoiding its spread. However, the similarities between common skin disorders and mpox and the need for professional diagnosis unavoidably impaired the diagnosis of earlier-stage mpox patients and contributed to mpox outbreak. To address the challenge, we proposed “Super Monitoring”, a real-time visualization technique employing artificial intelligence (AI) and Internet technology to diagnose earlier-stage mpox cheaply, conveniently, and quickly. Concretely, AI-mediated “Super Monitoring” (mpox-AISM) integrates deep learning models, data augmentation, self-supervised learning, and cloud services. According to publicly accessible datasets, mpox-AISM’s Precision, Recall, Specificity, and F1-score in diagnosing mpox reach 99.3%, 94.1%, 99.9%, and 96.6%, respectively, and it achieves 94.51% accuracy in diagnosing mpox, six like-mpox skin disorders, and normal skin. With the Internet and communication terminal, mpox-AISM has the potential to perform real-time and accurate diagnosis for earlier-stage mpox in real-world scenarios, thereby preventing mpox outbreak.
@@ 1 Introduction
Given the above issues, artificial intelligence (AI) offers a reliable solution [10]. Recently, deep learning (DL) models in AI have achieved significant success in various fields, including machine vision, medical imaging, and driverless vehicles. They have become sophisticated tools for addressing previously unknown and challenging problems. Moreover, AI-based innovations have reduced the cost of expensive equipment and professional knowledge required in the clinical diagnosis process, thereby helping overcome resource-constrained environments [11]. In this work, we employed the Internet, popular DL algorithms, data augmentation and a novel self-supervised learning (SSL) strategy, i.e. "A Simple Framework for Contrastive Learning of Visual Representations" (SimCLR) [12], to design a real-time and online strategy for diagnosing mpox. Our strategy considers the clinical characteristics of mpox rash and its high similarity to common skin diseases and helps diagnose earlier-stage mpox conveniently and swiftly. We validated the effectiveness of our strategy using publicly available datasets. Furthermore, we classified the mpox rash images into four grades (I, II, III, and Others) and two stages (earlier and later) based on dataset characteristics and the clinical evolution trend of the mpox rash. In detail, Grade-I means body parts are more prone to pox rash, i.e., the faces, necks, and hands [13, 14]. At the same time, Grade-II contains the arms and legs, which are not easily covered by clothing and other articles of daily use. We retested our strategy and found that it still performs well in diagnosing mpox images in different parts and stages. Particularly, to improve the reliability and security of our strategy, we also visualized and explained its decision-making process and results.
@@ 4 Limitations of study
AICOM-MP [36] 2024 SL 6124 labeled images mpox, non-mpox × 96.3% ×
MobileNetV3-large [37] 2024 SL 400 labeled images mpox, chickenpox, acne, normal × 88.2% ×
PoxNet22[38] 2023 SL 3192 labeled images mpox, non-mpox × 100.0% ×
DenseNet201[39] 2023 SL 1710 labeled images mpox, chickenpox, measles, normal × 97.6% ×
MobileNetV3-Small[40] 2023 SL 2056 labeled images mpox, non-mpox × 96.0% ×
ResNet18 [41] 2023 SL 3192 labeled images mpox, non-mpox × 99.5% ×
Vision Transformer [42] 2023 SL 3192 labeled images mpox, non-mpox × 94.7% ×
MonkeyNet [43] 2023 SL 8689 labeled images mpox, chickenpox, measles,normal × 93.2% ×
Xception+Denseet169 [44] 2022 SL 1754 labeled images mpox, chickenpox, measles, normal × 87.13% ×
MobileNetV2 [45] 2022 SL 3192 labeled images mpox, non-mpox × 91.1%
@@ 5 STAR METHODS
ResNet He et al. [53] %****␣templateArxiv.tex␣Line␣225␣****https://github.com/pytorch/vision/blob/main/torchvision/models/resnet.py